# Reconstructing MSM Sexual Networks to Guide PrEP Distribution Strategies for HIV Prevention

### v1

@@ Abstract
Men who have sex with men (MSM) remain disproportionately affected by HIV, yet optimizing the distribution of pre-exposure prophylaxis (PrEP) in this population remains a major public health challenge. Current PrEP eligibility guidelines and most modelling studies do not incorporate sociodemographic or network-level factors that shape transmission. We present a novel network reconstruction framework that generates MSM sexual contact networks from individual-level behavioral data, incorporating clustering and demographic assortativity by age, race, and sexual activity. Using data from 4,667 MSM participants, we reconstructed networks with varying topological properties and simulated HIV transmission over 50 years. Network structure strongly influenced outcomes: assortative by degree networks showed 18% lower equilibrium prevalence (63% vs 80% in race-assortative networks) due to hub isolation within communities. Targeted PrEP strategies based on degree or k-shell centrality achieved similar reductions with 20–40% coverage, matching random allocation at 60–80% coverage, particularly in assortative by age and race networks where hubs bridge demographic groups. Empirical PrEP distribution was suboptimal, underperforming by up to 30% compared with network-based strategies. Our findings demonstrate that integrating demographic mixing patterns into network reconstruction fundamentally alters optimal intervention design, offering a practical framework for improving HIV prevention in MSM populations where complete contact data are unavailable.
@@ Introduction
Men who have sex with men (MSM) are at disproportionate risk of HIV infection, accounting for 67% of sexually transmitted HIV cases in the United States [37]. This burden reflects not only biological susceptibility associated with the nature of sexual intercourse but also complex behavioral and structural factors, including partner concurrency, condomless sex, substance use, and elevated prevalence. While individual and behavioral factors contribute to HIV risk, effective prevention must also consider the network-level structures that sustain transmission.
@@
Pre-exposure prophylaxis (PrEP) has demonstrated high efficacy for HIV prevention in MSM across multiple trials and population studies [22, 30, 19]. Nevertheless, PrEP’s efficiency is highly dependent on the choice of the most appropriate eligibility guidelines, which depend on the estimation of the risk of HIV exposure based only on sex-related behaviors and vary between countries [49]. Current U.S. guidelines, for example, prioritize PrEP to individuals that reported in the last 6 months sexual intercourse with HIV-positive partners, other sexual transmitted diseases (STDs) and/or inconsistent condom use [45]. However, individual-level risk behaviors may not capture contact patterns and network-level transmission risk: an individual with moderate behavioral risk but high network centrality bridging multiple communities or connecting to many partners may pose greater transmission risk than someone with high-risk behaviors but few connections. Current PrEP allocation strategies ignore the structure of sexual networks despite their fundamental role in transmission dynamics [48, 20, 27, 38].
@@
Network-based models studies have provided insights of optimized strategies to allocate PrEP among MSM, suggesting that targeting highly connected individuals can improve intervention efficiency [9, 15, 36, 6, 44]. However, existing approaches face critical limitations: they either require complete contact data, which is extremely difficult to collect for stigmatized populations, or rely on simplified topologies that fail to capture real-world network features. Moreover, MSM sexual networks exhibit distinctive structural properties that fundamentally influence transmission but remain absent from intervention planning frameworks. In particular, studies have found that people tend to connect or choose potential sexual partners within the same age or racial groups [1, 3]. Members of the same sexual network often exhibit similar norms, attitudes, and HIV risk behavior levels [3]. Network-tailored public health policies could substantially reduce transmission chains among high-risk individuals by identifying and prioritizing those whose network position, along with individual behavior, makes them critical to epidemic dynamics.
@@
In this work, we address these gaps by leveraging data from 4,667 MSM participants in the ARTnet study [50] to: (i) develop and validate degree-preserving network reconstruction algorithms that uniquely incorporate demographic assortativity by age, race, and sexual activity, thus leading to networks with realistic structures; (ii) simulate HIV transmission dynamics across networks with different topological properties; (iii) compare PrEP allocation strategies based on network centrality (degree-based and k-shell decomposition) versus random distribution; and (iv) benchmark these approaches against empirically observed PrEP uptake patterns (see Figure 1).
@@
Our findings demonstrate that network structure significantly shapes both HIV transmission and PrEP effectiveness, with important implications for public health policy. Targeted strategies, focusing on centrality measures, consistently outperform current practices of PrEP deployment, particularly in networks where highly connected individuals bridge demographic groups. These results provide a practical framework for optimizing PrEP distribution when complete network data is unavailable, highlighting critical opportunities to improve HIV prevention in MSM communities.
@@ Results
We analyzed data from the ARTnet study [50], an anonymous cross-sectional Internet-based survey that collected information on sociodemographic characteristics, HIV-related risk behavior, testing, and the use of prevention strategies or services of 4904 MSM in the United States, between 2017 and 2019. To take part in the study, participants had to identify as a male cisgender and biologically male, have a history of sexual intercourse with other men, and be between 15 and 65 years old. After preprocessing the data (see Methods Section A), we obtained a cohort of 4667 MSM from which we extracted age, HIV status, number of sexual partnerships, race, PrEP status, and testing frequency [50].
By using this cohort, we were able to recreate a synthetic population that expresses real demographic and behavioral characteristics sampled from a MSM community in the United States. The median (P25-P75) number of partnerships and testing frequency was, respectively, 4 (2-10) and 2 (1-4). The most prevalent race group was white (72%), followed by hispanic (14%), other (9%), and black (5%). The median (P25-P75) age was 32 (24-49). Distributions of number of contacts, age, race, and testing frequency are given in the Supplementary Information Fig. S3.
@@
The ARTNet dataset only contains individual-level information without contact patterns structure. We overcome this limitation by proposing a novel reconstruction framework to generate synthetic sexual networks based on sociodemographic data. First, we implemented the Havel-Hakimi algorithm [24, 23] to generate networks that preserve the observed degree sequence ($V=4,667$ nodes, $E=25,107$ edges), followed by a Markov Chain Monte Carlo (MCMC)-based edge swapping to ensure proper randomization (see Methods Section B, Supplementary Information Algorithm 1 and 2). To ensure statistical robustness, we ran 5 independent Markov chains, each initialized with 8 different networks generated by the Havel-Hakimi algorithm. Each chain ran until convergence was reached. This ensemble of 40 random networks was used as a baseline to systematically enhance, independently, specific structural properties while maintaining degree distributions constant.
While previous studies have generated networks with clustering and degree assortativity [7, 33, 46], our key contribution lies in incorporating demographic assortativity, more specifically by age and race, into network reconstruction. Real sexual contact networks reflect not only who connects to whom (how many partners – degree), but crucially the tendency of individuals to connect to those who share similar characteristics.[3]. We independently implement and evaluate how different forms of assortativity influence epidemic dynamics based on four network properties: (i) clustering coefficient, (ii) degree assortativity, (iii) assortativity by age, and (iv) assortativity by race. This allows us to create networks that range from random mixing to highly segregated demographic communities (see Methods Section B and Supplementary Information Section 4).
@@
To understand how demographic mixing patterns influence network structure beyond simple degree distributions, we compared networks reconstructed using traditional methods (clustering and degree assortativity) with our proposed demographic-based approach (age and race assortativity). We characterized such networks across multiple scales, from global topology to mesoscale community organization.
@@
At the global-scale level, the network topology of each generated network was characterized by computing the largest component (LC), average path length ($\langle l\rangle$), diameter ($D$), and modularity ($Q$) as we increased the property under study (see Supplementary Information Fig. S5 and S7 (a)). Table 1 shows the metrics of the most enhanced clustered and assortative by degree, age and race networks, averaged across all network realizations. We observe that assortative by degree and clustered networks exhibit the most pronounced structural changes, with the longest average path lengths ($\langle l\rangle=6.23\pm 4.85\times 10^{-2}$ and $4.84\pm 2.01\times 10^{-2}$, respectively) and bigger diameters ($D=21.68\pm 1.794$ and $12.25\pm 9.080\times 10^{-1}$), while maintaining substantial connectivity despite increased fragmentation. Modularity remained similar across clustered ($Q=0.558\pm 4.92\times 10^{-3}$), assortative by degree ($Q=0.573\pm 1.00\times 10^{-2}$), and assortative by age ($Q=0.600\pm 9.49\times 10^{-3}$) networks. In contrast, assortative by race networks showed minimal structural deviation from the random networks, presenting the lowest fragmentation ($DC=11.95\pm 3.794$) and modularity ($Q=0.373\pm 3.21\times 10^{-3}$), suggesting weaker racial homophily in sexual contact patterns.
@@
Table 1: Structural metrics of networks with maximally enhanced properties. From an ensemble of 40 baseline networks (generated via 5 MCMC chains × 8 initial havel-hakimi network configurations), we enhanced each property through edge rewiring and selected those achieving maximum enhancement. Values show the mean and standard deviation of the final enhanced property under study ($P_{f}$), the largest component (LC), number of disconnected components (DC), diameter ($D$), average path length ($\langle l\rangle$) and modularity ($Q$).
$P_{f}$ LC DC $D$ $\langle l\rangle$ Q
Clustering $0.416\pm 1.64\times 10^{-3}$ $4608\pm 13.06$ $29.35\pm 6.002$ $12.25\pm 9.080\times 10^{-1}$ $4.84\pm 2.01\times 10^{-2}$ $0.558\pm 4.92\times 10^{-3}$
Assortative By Degree $0.486\pm 5.64\times 10^{-4}$ $3931\pm 24.97$ $321.7\pm 8.429$ $21.68\pm 1.794$ $6.23\pm 4.85\times 10^{-2}$ $0.573\pm 1.00\times 10^{-2}$
Assortative By Age $0.974\pm 2.49\times 10^{-4}$ $4642\pm 6.811$ $3.45\pm 3.193 11$$1.50\pm 5.478\times 10^{-1}$ $4.42\pm 2.11\times 10^{-2}$ $0.600\pm 9.49\times 10^{-3}$
Assortative By Race $0.793\pm 2.80\times 10^{-3}$ $4645\pm 7.641$ $11.95\pm 3.794$ $8.350\pm 5.268\times 10^{-1}$ $3.54\pm 6.09\times 10^{-3}$ $0.373\pm 3.21\times 10^{-3}$
At the meso-scale level, we characterized the community structure of each type of network, using a single stochastic realization, given the minimal standard deviation in global network metrics across trials (see Table 1). Community structure analysis using the Nested Stochastic Block Model (NSBM)[42] – chosen due to its robustness (see Supplementary Information, Fig. S6 and Section 5) – revealed significant differences despite similar modularity values. The number of optimal communities varied significantly: clustered networks fragmented into 56 communities, while assortative by race networks formed only 6 (see Figures 2 and 3, and Supplementary Information, Fig. S8). Community size distribution was quite distinct across different networks. In clustered and assortative by degree networks, nodes were concentrated in single large communities (Communities 11 and 14, respectively), while other communities were significantly smaller than the average community size. Assortative by age networks showed the most homogeneous community size distribution, while assortative by race networks demonstrated the greatest concentration of nodes into few specific communities, with Community 1 containing approximately 70% of all nodes (see Figure 2).
@@
Another distinct finding was the heterogeneous distribution of highly connected individuals across the networks we generated (see Figure 3). In assortative by degree and clustered networks, hubs concentrated within communities with few nodes (Community 0 and 1, and Community 18, 41 and 54, respectively), creating isolated transmission chains. In contrast, the largest communities in these networks (Communities 14 and 11, respectively) showed the lowest median degrees. On the other hand, assortative by age and race networks presented hubs distributed more evenly across communities and typically resided in lower-degree communities. This pattern was particularly pronounced in assortative by race networks where despite Community 2’s high median degree, most hubs concentrated in the dominant Community 1 (see Figure 2 (d)).
@@
The proposed framework to tune demographic mixing patterns while preserving degree distributions represents a methodological advance that enables us to isolate how different forms of assortativity influence epidemic dynamics. To assess how global and mesoscopic topological properties influence HIV spreading among MSM contact networks, we simulated HIV transmission dynamics using a discrete-time stochastic agent-based model over 50 years in the largest connected component of each type of network. In our model, each node is represented by an individual which is assumed to be in one of the following states: susceptible, infected or on treatment. Transmission probability was highly dependent on the number of infected neighbors, sexual role, condom use (30% of coverage [29] and 0.70 effectiveness [31]), and PrEP status. Individuals who are on treatment were assumed virally suppressed and non-infectious [13] (see Methods Section C for a full description of the agent-based model implementation). Simulations were initialized with approximately 8.6% people living with HIV (PLHIV) across all networks matching empirical data, with no PrEP intervention at baseline.
Epidemic trajectories were quantified by measuring the proportion of PLHIV, accounting for both infected and on treatment individuals. Baseline simulations revealed substantial variation in equilibrium prevalence: assortative by degree networks showed the lowest burden (63.2%, IQR: 62.8–63.7%) and fastest convergence, while assortative by age and race networks reached the highest prevalence (78.1%, IQR: 77.5–78.4% and 80.1%, IQR: 79.8–80.4, respectively) (see Supplemmentary Information, Fig. S7 (b)). These differences directly reflect the structural properties characterized above. Assortative by degree networks’ increased fragmentation, longer path lengths, and hub isolation created natural transmission barriers, while assortative by race networks’ lower modularity facilitated broader spread, a finding with important implications for targeted intervention design.
@@
The aforementioned baseline dynamics were used as benchmarks to evaluate PrEP allocation strategies. We systematically varied PrEP individual adherence (0.23, 0.80, 0.95) and coverage, the fraction of susceptible individuals receiving PrEP (0.05-0.80). The adherence level is proportional to PrEP dosage uptake, such that higher adherence corresponds to lower HIV susceptibility. In the beginning of the dynamics, for different coverage levels, a fraction of susceptible nodes is chosen to be on PrEP. A person is eligible to be chosen for PrEP uptake if they are HIV-negative [32]. The selected individuals are assumed to maintain the adoption of the respective prevention strategies during the entire simulation. Additionally, we compared three allocation approaches: random distribution, degree centrality targeting, and k-shell decomposition targeting (see Methods Section E). Strategy effectiveness was quantified as the reduction in cumulative final size of PLHIV relative to baseline ($F_{s}$$F_{0}$) over 50-year simulations (see Figure 4).
@@
At maximum adherence (95%), targeted strategies substantially outperformed random allocation across all network types, requiring lower coverage to achieve equivalent reductions in PLHIV. Assortative by age and race networks showed the greatest sensitivity to targeted interventions: degree-based targeting reduced prevalence to $F_{s}/F_{0}\approx 0.25$ at only 40% coverage in age-assortative networks, while random distribution required 80% coverage to achieve similar reductions (see Figure 4 (i) and (l), respectively). Degree centrality consistently outperformed k-shell targeting in assortative by age networks at lower coverage levels ($F_{s}/F_{0}\approx 0.55$ vs. $0.70$ at 20% coverage), with performance converging at higher coverage (see Figure 4 (h) and (i)). Such divergence can be explained by the lower correlations between degree and k-shell metrics in assortative by age networks (see Supplementary Information, Fig. S10 (c)).
We also compared empirical PrEP uptake distribution, retrieved from the data, against simulated strategies at maximum adherence (95%), mapping coverage-level observed in the data to those in the simulation (see Figure 5).
@@
Since network fragmentation varied across topologies, we adjusted coverage to reflect the proportion of PrEP users in each network’s largest component. Consequently, in maximally clustered networks, $13.8\%$ of nodes were allocated PrEP, while for networks with maximal assortativity by degree, age and race this percentage was $15.8\%$, $13.8\%$ and $13.7\%$, respectively. Empirical distribution matched or exceeded simulated strategies in clustered and assortative by degree networks, where PrEP coverage exceeded 20% among medium to high-degree individuals (see Supplementary Information, Fig. S9 (b)). This suggests current uptake includes network-influential individuals in these topologies. However, in networks where assortativity is driven by age or race, we found that the empirical PrEP distribution is suboptimal (see Figure 5 (c) and (d)). In this case, both degree and k-shell targeting substantially outperformed empirical PrEP distribution, particularly in assortative by race networks, where coverage averaged only $\sim$10% within demographic groups (see Supplementary Information, Fig. S9 (a)). In these networks, hubs bridge demographic communities, yet our results reflect that in these topologies crucial individuals are not on PrEP. This may be due to current eligibility criteria being focused on individual risk behaviors, failing to identify these more nuanced transmission-critical individuals. These findings reveal substantial opportunities for optimization through network-informed allocation strategies that explicitly prioritize individuals bridging demographic groups or occupying central network positions.

### v2

@@ Abstract
Men who have sex with men (MSM) remain disproportionately affected by HIV, yet optimizing Pre-exposure Prophylaxis (PrEP) distribution remains a public health challenge. Current guidelines and most modelling studies do not incorporate sociodemographic or network-level factors that shape transmission. While network reconstruction from egocentric data has been studied, the relative importance of demographic mixing dimensions remains uncertain. Using data from 4,667 MSM participants, we show that uncertainty in network reconstruction from egocentric survey data — specifically whether assortativity by age or race is incorporated — affects simulated HIV prevalence under the same observed PrEP uptake. We simulate HIV transmission over 50 years across this structural space and evaluate whether empirically observed uptake reaches transmission-critical network positions. Network structure strongly influences outcomes: assortative by degree networks show 17% lower equilibrium prevalence due to hub isolation within communities. Targeted PrEP strategies based on degree or k-shell centrality achieved the highest prevalence reductions, particularly in assortative by age and race networks where hubs bridge demographic groups. PrEP uptake from data is suboptimal in assortative by age and race networks, underperforming compared with network-based strategies. Results demonstrate that uncertainty in network reconstruction affects intervention design and highlight the need for robust prevention strategies under structural ambiguity.
@@ Introduction
Men who have sex with men (MSM) are at a disproportionate risk of HIV infection, accounting for 67% of sexually transmitted HIV cases in the United States [40]. This burden reflects not only biological susceptibility associated with the nature of sexual intercourse but also complex behavioral and structural factors, including partner concurrency, condomless sex, substance use, and elevated HIV prevalence. While individual and behavioral factors contribute to HIV risk, effective prevention must also consider the network-level structures that sustain transmission.
@@
Pre-exposure prophylaxis (PrEP) has demonstrated high efficacy for HIV prevention in MSM across multiple trials and population studies [24, 33, 21]. Nevertheless, PrEP’s efficiency is highly dependent on the choice of the most appropriate eligibility guidelines. These depend on the estimation of the risk of exposure to HIV based solely on sex-related behaviors and vary between countries [53]. For example, current U.S. guidelines prioritize PrEP for individuals who, in the last six months, have reported sexual intercourse with HIV-positive partners, other sexually transmitted diseases (STDs), and/or inconsistent condom use[49]. However, individual-level risk behaviors may not capture contact patterns and network-level transmission risk. An individual with moderate behavioral risk but high network centrality, bridging multiple communities or connecting to many partners, may pose greater transmission risk than someone with high-risk behaviors but few connections. Because sexual network structure is not directly observable in routine surveillance, PrEP allocation is designed without knowledge of individuals’ positions in the contact network, leaving open the question of whether current uptake reaches the most transmission-critical individuals [52, 22, 30, 41]. Nonetheless, previous work has shown that risk-based targeting is not always the most effective in practice, particularly in heterogeneous efficacy scenarios [48].
@@
Network-based models studies have provided valuable insights on optimized strategies to allocate PrEP among MSM, suggesting that targeting highly connected individuals can improve intervention efficiency [10, 17, 39, 6, 47]. However, existing approaches face critical limitations: they either require complete contact data, which is extremely difficult to collect for stigmatized populations, or rely on simplified topologies that fail to capture real-world network features. Moreover, MSM sexual networks exhibit distinctive structural properties that fundamentally influence transmission but remain absent from intervention planning frameworks. In particular, studies have found that people tend to connect or choose potential sexual partners within the same age or racial groups [1, 3]. Members of the same sexual network often exhibit similar norms, attitudes, and HIV risk behavior levels [3]. Network-tailored public health policies could substantially reduce transmission chains among high-risk individuals by identifying and prioritizing those whose position in a given network structure, along with individual behavior, makes them critical to epidemic dynamics.
@@
In this work, we leverage data from 4,667 MSM participants in the ARTnet study [54], of whom 13.67% were on PrEP at the time of data collection, to: (i) develop and validate degree-preserving network reconstruction algorithms that uniquely incorporate demographic assortativity by age, race, and sexual activity, thus leading to networks with realistic structures; (ii) simulate HIV transmission dynamics across networks with different topological properties; (iii) compare PrEP allocation strategies based on network centrality — degree-based and k-shell decomposition — against random distribution across this structural space, quantifying how network topology modulates intervention effectiveness; and (iv) assess whether observed PrEP uptake from the data takes transmission-critical network positions, or whether there is an under-representation of individuals occupying central or bridging roles across different network reconstructions (see Figure 1).
@@
Our findings show that networks reconstructed assuming assortative mixing with respect to different attributes, including age, race and activity level, can differ significantly, leading to different HIV transmission patterns and PrEP effectiveness, with important implications for public health policy. Targeted strategies, focusing on centrality measures, consistently outperform observed PrEP uptake from the data, particularly in networks where highly connected individuals bridge demographic groups. These results provide evidence of the need for a practical framework when complete network data is unavailable, highlighting opportunities to suggest improvements in HIV prevention and optimized PrEP distribution in MSM communities.
@@ Results
We use data from the ARTnet study [54], an anonymous cross-sectional Internet-based survey that collected information on sociodemographic characteristics, HIV-related risk behavior, testing, and the use of prevention strategies or services of 4904 MSM in the United States, between 2017 and 2019. To take part in the study, participants had to identify as a male cisgender and biologically male, have a history of sexual intercourse with other men, and be between 15 and 65 years old. After preprocessing the data (see Methods Section A), we obtained a cohort of 4667 MSM from which we extracted age, HIV status, number of sexual partnerships, race, PrEP status, and testing frequency [54].
With this cohort, we create a synthetic population that expresses real demographic and behavioral characteristics sampled from a MSM community in the United States. The median (P25-P75) number of partnerships and the testing frequency are, respectively, 4 (2-10) and 2 (1-4). The most prevalent race group is white (72%), followed by hispanic (14%), other (9%), and black (5%). The median (P25-P75) age is 32 (24-49). The proportion of participants reporting being currently on PrEP was 13.67%, consistent with 2017 CDC estimates for MSM [14]. Complete distributions of number of contacts, age, race, and testing frequency are given in the Supplementary Information Fig. S3.
@@
The ARTNet dataset contains only individual-level information without contact patterns structure. To overcome this limitation, we propose a reconstruction framework that generates synthetic sexual networks from sociodemographic data and explores how network structure changes when assortativity is imposed with respect to different demographic attributes individually. First, we implemented the Havel-Hakimi algorithm [26, 25] to generate networks that preserve the observed degree sequence ($V=4,667$ nodes, $E=25,107$ edges), followed by a Markov Chain Monte Carlo (MCMC)-based edge swapping to ensure proper randomization (see Methods Section B, Supplementary Information Algorithm 1 and 2). To ensure statistical robustness, we run 10 independent Markov chains, each initialized with 10 different networks generated by the Havel-Hakimi algorithm. Each chain runs until convergence is reached. This ensemble of 100 random networks serves as a baseline for independently modifying specific structural properties while keeping the degree distribution fixed. Real sexual contact networks reflect not only who connects to whom (how many partners – degree), but most importantly how individuals tend to connect to those who share similar characteristics.[3]. Thus, beyond common approaches that generate networks with clustering and degree assortativity [8, 36, 50], we incorporate assortative mixing by demographic attributes, specifically age and race, into network reconstruction, revealing their distinct structural effects on HIV transmission dynamics. By reconstructing networks based on these four network properties, we create networks that range from random mixing to highly segregated demographic communities (see Methods Section B and Supplementary Information Section 4).
@@
To examine how assortativity by different attributes shapes network structure beyond degree distributions, we compare networks reconstructed using traditional methods (clustering and degree assortativity) with our demographic-based approach (age and race assortativity). We characterize such networks across multiple scales, from global topology to mesoscale community organization.
@@
At the global-scale level, the network topology of each generated network is characterized by computing the largest component (LC), average path length ($\langle l\rangle$), diameter ($D$), and modularity ($Q$) while systematically varying the structural property of interest (see Supplementary Information Fig. S5 and S6 (a)). Modularity was estimated by partitioning the network using the Louvain algorithm, as implemented in the python-louvain package [7]. All remaining global metrics were computed using NetworkX. Table 1 shows the metrics of networks with maximal clustering and maximal assortativity by degree, age, and race, averaged across all network realizations. We observe that assortative by degree and clustered networks exhibit the most pronounced structural changes, with the longest average path lengths ($\langle l\rangle=6.23\pm 0.0526$ and $4.84\pm 0.0192$, respectively) and bigger diameters ($D=21.6\pm 1.84$ and $12.3\pm 0.853$), while maintaining substantial connectivity despite increased fragmentation. Modularity remains similar across clustered ($Q=0.558\pm 0.00435$), assortative by degree ($Q=0.573\pm 0.00902$), and assortative by age ($Q=0.601\pm 0.00782$) networks. In contrast, assortative by race networks show minimal structural deviation from the random networks, presenting the lowest fragmentation ($DC=11.6\pm 3.33$) and modularity ($Q=0.373\pm 0.00353$), suggesting weaker racial homophily in sexual contact patterns.
@@
Table 1: Structural metrics of networks with maximally enhanced properties. From an ensemble of 100 baseline networks (generated via 10 MCMC chains × 10 initial Havel–Hakimi network configurations), we optimized each property through edge rewiring and selected those networks attaining the largest increase in the target metric. Values show the mean and standard deviation of the final property value ($P_{f}$), the largest component (LC), number of disconnected components (DC), diameter ($D$), average path length ($\langle l\rangle$), and modularity ($Q$).
$P_{f}$ LC DC D <l> Q
Clustering 0.416 ± 0.00154 4.61e+03 ± 11.6 28.8 ± 5.31 12.3 ± 0.853 4.84 ± 0.0192 0.558 ± 0.00435
Assortative By Degree 0.486 ± 0.000574 3.93e+03 ± 23.2 323 ± 7.65 21.6 ± 1.84 6.23 ± 0.0526 0.572 ± 0.00902
Assortative By Age 0.974 ± 0.000248 4.64e+03 ± 7.3 13.8 ± 3.53 11.5 ± 0.623 4.42 ± 0.0244 0.601 ± 0.00782
Assortative By Race 0.793 ± 0.00271 4.65e+03 ± 6.74 11.6 ± 3.33 8.33 ± 0.53 3.54 ± 0.00606 0.373 ± 0.00353
@@
At the meso-scale level, we performed community structure analysis using the Nested Stochastic Block Model (NSBM)[45]. Given the stochastic nature of network generation, we assessed community structure consistency across realizations for each network type using Kolmogorov-Smirnov (KS) tests to compare pairwise community size distributions. Under the null hypothesis that these distributions are identical, most within type comparisons failed to reject $H_{0}$ ($p>0.05$) (see Supplementary Information, Fig. S7 and S8 and Section 5), indicating strong meso-scale structure similarity among networks generated under the same assortativity and clustering parameter constraints. Therefore, one stochastic trial was chosen to represent the ensemble of networks generated for each network type, from which the community structure was characterized (see Figure 2 and 3). While within each network type, there was consistency in the community size distributions, the same distributions revealed great differences when compared across different network types, despite similar modularity values (see Table 1).
The number of optimal communities varied significantly: clustered networks fragmented into 56 communities, while assortative by race networks formed only 6 (see Figures 2 and 3, and Supplementary Information, Fig. S9). Community size distribution was quite distinct across different networks. In clustered and assortative by degree networks, nodes were concentrated in single large communities (Communities 11 and 14, respectively), while other communities were much smaller than the average community size. Assortative by age networks showed the most homogeneous community size distribution, while assortative by race networks demonstrated the greatest concentration of nodes into few specific communities, with Community 1 containing approximately 70% of all nodes (see Figure 2).
@@
We also observed a significantly heterogeneous distribution of highly connected individuals across the generated networks (see Figure 3). In assortative by degree and clustered networks, hubs concentrated within communities with few nodes (Community 0 and 1, and Community 18, 41 and 54, respectively), creating isolated transmission chains. In contrast, the largest communities in these networks (Communities 14 and 11, respectively) showed the lowest median degrees. On the other hand, assortative by age and race networks presented hubs distributed more evenly across communities and typically resided in lower-degree communities. This pattern was particularly pronounced in assortative by race networks where despite Community 2 high median degree, most hubs concentrated in the dominant Community 1 (see Figure 2 (d)).
@@
The proposed framework to explore extremes of demographic mixing patterns while preserving degree distributions represents a methodological advance that allows us to isolate how different forms of assortativity influence epidemic dynamics. To assess how mesoscopic and global topological properties influence HIV spreading among these sexual contact networks, we simulated HIV transmission dynamics using a discrete-time stochastic agent-based model over 50 years in the largest connected component of each network type (see Methods Section C). Consistent with patterns observed in community structure, the largest component size was stable across network realizations for each type. (see Table 1 and Supplementary Information, Figure S6 (a)). Accordingly, we used the same network analyzed for community structure to perform our epidemic simulations (see Figure 2 and 3). In our model, each node is represented by an individual which is assumed to be in one of the following states: susceptible, infected or on treatment. Transmission probability was highly dependent on the number of infected neighbors, sexual role, condom use (30% of coverage [32] and 0.70 effectiveness [34]), and PrEP status. Individuals who are on treatment were assumed virally suppressed and non-infectious [15] (see Methods Section C for a full description of the agent-based model implementation). Simulations were initialized with approximately 8.6% people living with HIV (PLHIV) in the largest component of all networks, with no PrEP intervention at baseline.
Epidemic trajectories were quantified by measuring the proportion of PLHIV, accounting for both infected and on treatment individuals. Baseline simulations revealed substantial variation in equilibrium prevalence: assortative by degree networks showed the lowest burden (63.2%, IQR: 62.8–63.7%) and fastest convergence, while assortative by age and race networks reached the highest prevalence (78.1%, IQR: 77.5–78.4% and 80.1%, IQR: 79.8–80.4, respectively) (see Supplementary Information, Fig. S6 (b)). These differences directly reflect the structural properties characterized above. Assortative by degree networks’ increased fragmentation, longer path lengths, and hub isolation created natural transmission barriers, while assortative by race networks’ lower modularity facilitated broader spread, a finding with important implications for planning targeted intervention design.
@@
The aforementioned baseline dynamics was used as a benchmark to evaluate PrEP allocation strategies. We systematically varied PrEP individual adherence (0.23, 0.80, 0.95) and coverage (0.05-0.80). The adherence level is proportional to PrEP dosage uptake, such that higher adherence corresponds to lower HIV susceptibility. PrEP coverage corresponds to the fraction of susceptible individuals receiving PrEP. At the beginning of each simulation and for each coverage level, a corresponding fraction of susceptible nodes is selected to be on PrEP. An individual is eligible to be chosen for PrEP uptake if they are HIV-negative [35]. The selected individuals are assumed to maintain the adoption of the respective prevention strategies during the entire simulation. Additionally, we compared three allocation approaches: random distribution, degree centrality targeting, and k-shell decomposition targeting (see Methods Section E). Strategy effectiveness was quantified as the reduction in cumulative final size of PLHIV relative to baseline ($F_{s}$$F_{0}$) over 50-year simulations (see Figure 4).
@@
At maximum adherence (95%), targeted strategies substantially outperformed random allocation across all network types, requiring lower coverage to achieve equivalent reductions in PLHIV. Assortative by age and race networks showed the greatest sensitivity to targeted interventions: degree-based targeting reduced prevalence to $F_{s}/F_{0}\approx 0.30$ at only 40% coverage in age-assortative networks, while random distribution required 80% coverage to achieve similar reductions (see Figure 4 (i) and (l), respectively). Degree centrality consistently outperformed k-shell targeting in assortative by age networks at lower coverage levels ($F_{s}/F_{0}\approx 0.55$ vs. $0.70$ at 20% coverage), with performance converging at higher coverage (see Figure 4 (h) and (i)). Such divergence can be explained by the lower correlations between degree and k-shell metrics in assortative by age networks (see Supplementary Information, Fig. S11 (c)).
Given that PrEP empirical data matches national reported PrEP uptake levels among MSM [14], we also compared PrEP uptake retrieved from the data with the simulated strategies at maximum adherence (95%), mapping coverage-level observed in the data with those in the simulation (see Figure 5).
@@
Since network fragmentation varied across topologies, we adjusted coverage to reflect the proportion of PrEP users in each network’s largest component. Consequently, in maximally clustered networks, $13.8\%$ of nodes were allocated PrEP, while for networks with maximal assortativity by degree, age and race this percentage was $15.8\%$, $13.8\%$ and $13.7\%$, respectively. Uptake from the data matched or exceeded simulated strategies in clustered and assortative by degree networks, where PrEP coverage exceeded 20% among medium to high-degree individuals (see Supplementary Information, Fig. S10 (b)). This suggests that current uptake covers network-influential individuals in these topologies. However, in networks where assortativity is driven by age or race, the observed PrEP uptake does not align with network-central positions, suggesting that if real MSM networks exhibit this structure (see Figure 5 (c) and (d)), there may be scope for complementing current broad-access strategies with outreach focused on individuals bridging age or race clusters. In this case, both degree and k-shell targeting substantially outperformed the observed PrEP uptake, particularly in assortative by race networks, where coverage averaged only $\sim$10% within demographic groups (see Supplementary Information, Fig. S10 (a)). In these networks, hubs bridge demographic communities, yet our results suggest that in these topologies crucial individuals may not be on PrEP. Given that current eligibility criteria are focused on individual risk behaviors, we show that there is an advantage in also identifying these more nuanced transmission-critical individuals. In general, these findings reveal substantial opportunities for optimization through network-informed allocation strategies that explicitly address individuals bridging demographic groups or occupying central network positions, in addition to current guidelines. Moreover, in scenarios where there could be limited resources and PrEP distribution, it is crucial to understand how to optimally distribute it.